OpenStudy (anonymous):

the world's population grow at the rate of 2% per year. if it assumed that the population growth is exponential, then the population t year from now will be given by a function of the form p(t)=p0 e^0.02t, where p0 is the current population. Assuming that this model of population growth is correct, how long will it take for the world’s population to double?

OpenStudy (anonymous):

Solve for t: e^(0.02t)=2

OpenStudy (anonymous):

so... e^(0.02t)=2 0.02t=ln2 t=ln2/0.02 so t = is my answer?

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Anytime :)